OpenStudy (lgbasallote):

How many possible 4 letter words can be made from the word "LGBASALLOTE"?

OpenStudy (baldymcgee6):

real words?

OpenStudy (lgbasallote):

just possible 4 letter arrangements

OpenStudy (lgbasallote):

and it's actually much harder than it seems...

OpenStudy (baldymcgee6):

\[\frac{ 11! }{ 3! * 2! }\]

OpenStudy (lgbasallote):

that's the permutation for the whole word...

OpenStudy (lgbasallote):

i would assume cases are needed ehre

OpenStudy (anonymous):

First you use \[ \left( \begin{array}{c} 11 \\ 4 \end{array} \right) \]To get the total number of words that you could get if there were no indistinguishable letters. Then you have to figure out a way to get rid of double counting.

OpenStudy (lgbasallote):

hmm...i get that first part....

OpenStudy (anonymous):

similar question

OpenStudy (lgbasallote):

that's an easy one...it just has 2 repeating characters.

OpenStudy (lgbasallote):

but what if there are more?

OpenStudy (anonymous):

take different cases..

OpenStudy (anonymous):

For every word that uses two A, it won't be double counted, but if it uses one A, it will be double counted.

OpenStudy (anonymous):

If it has 1 L, then it will be tripled counted, if it has two Ls, it will be counted 3 chose 2 times, if it has 3 L it won't be double counted

OpenStudy (anonymous):

There isn't necessarily a nice looking formula for this.

OpenStudy (anonymous):

\[ \left( \begin{array}{c} 11 \\ 4 \end{array} \right) - \left( \begin{array}{c} 3 \\ 1 \end{array} \right) \left( \begin{array}{c} 8-3 \\ 3 \end{array} \right) - \left( \begin{array}{c} 3 \\ 2 \end{array} \right) \left( \begin{array}{c} 8-3 \\ 2 \end{array} \right) - \left( \begin{array}{c} 3 \\ 3 \end{array} \right) \left( \begin{array}{c} 8-3 \\ 1 \end{array} \right) \]\[- \left( \begin{array}{c} 2 \\ 1 \end{array} \right) \left( \begin{array}{c} 8-2 \\ 3 \end{array} \right) - \left( \begin{array}{c} 2 \\ 2 \end{array} \right) \left( \begin{array}{c} 8-2 \\ 2 \end{array} \right) \]Something like this...

OpenStudy (anonymous):

Though I might be overlooking something.

OpenStudy (lgbasallote):

....wow.....

OpenStudy (anonymous):

Those 8s should be 11s

OpenStudy (lgbasallote):

why subtraction?

OpenStudy (anonymous):

It's my best attempt at trying to find any overlap.

terenzreignz (terenzreignz):

LGBASALLOTE There are 3 L's and 2 A's, nothing else is repeated. Getting rid of the repeated letters for now... LGBASOTE For 4 letter words with no repeated letters, there are 8P4 possibilities. Suppose there are 2 L's and no other repetitions There are 6 (4C2) ways to scatter 2 L's around a 4 letter word times the number of ways the 7 non-L letters could be scattered around the two remaining "slots" So that's 6 x 7P2 Suppose there are 2 A's and no other repetitions There are 6 (4C2) ways to scatter 2 A's around a 4 letter word times the number of ways the 7 non-A letters could be scattered around the two remaining "slots" so that's 6 x 7P2 Suppose there are 3 L's (of course, no other repetition may occur, as there is only one more letter) There are 4 (4C3) ways to scatter 3 L's around a 4 letter word times the number of ways the 7 non-L letters could be placed in the remaining "slot" so that's 4 x 7P1 = 4x7 = 28. Suppose there are 2 L's and 2 A's Then the number of ways they can be scattered is equal to the number of ways two L's can be scattered around a 4 letter word so that's 6. TOTAL 8P4 + (6)(7P2) + (6)(7P2) + 28 + 6

OpenStudy (lgbasallote):

why 8P4?

OpenStudy (lgbasallote):

shouldn't it be 6P4?

terenzreignz (terenzreignz):

Because there are 8 distinct letters?

OpenStudy (lgbasallote):

so you didn't consider the arrangement where there are no L or A?

terenzreignz (terenzreignz):

No I didn't... Thanks for the heads-up Recalculating :D

OpenStudy (lgbasallote):

same thing for @wio why 11C4?

terenzreignz (terenzreignz):

Wait, 8P4 already does account for the possibilities that there may be no L or A.

OpenStudy (lgbasallote):

it does?

terenzreignz (terenzreignz):

I think so. You simply cannot use all 8 letters in a 4-letter word, after all.

OpenStudy (lgbasallote):

hmm...i don't seem to find any mistake in that solution....

OpenStudy (lgbasallote):

except for that last 6....

OpenStudy (lgbasallote):

shouldn't it be \[\huge _3 C _ 2 \times _2C_2 \times \frac{4!}{2!2!}\]

OpenStudy (lgbasallote):

making it equal to 18....

terenzreignz (terenzreignz):

The last 6? When the 4 letter word consists purely of 2 L's and 2 A's?

OpenStudy (lgbasallote):

yes

terenzreignz (terenzreignz):

Well, it's only 6, after all, brute force gives us LLAA LALA LAAL ALLA ALAL AALL

OpenStudy (lgbasallote):

yes...that's the possible way of arranging...but what about the possible ways of picking the L?

OpenStudy (lgbasallote):

doesn't that make a difference?

terenzreignz (terenzreignz):

It shouldn't, unless you mean to say, for instance LGBA can be counted three times depending on which "L" I used... If it does, then it basically means all my calculations were near-useless...

OpenStudy (lgbasallote):

.....i suppose i can't prove you wrong then....

terenzreignz (terenzreignz):

^Which doesn't necessarily mean I'm right LOL

OpenStudy (anonymous):

8C4 *4! + 8C3*4!/2! + 8C2 * 4!/3! + 8C3*4!/2! +8C2 *4!/2!2!